# A Survey of Pathology Foundation Model: Progress and Future Directions

### v1

@@ Abstract
Computational pathology, analyzing whole slide images for automated cancer diagnosis, relies on the multiple instance learning framework where performance heavily depends on the feature extractor and aggregator. Recent Pathology Foundation Models (PFMs), pretrained on large-scale histopathology data, have significantly enhanced capabilities of extractors and aggregators but lack systematic analysis frameworks. This survey presents a hierarchical taxonomy organizing PFMs through a top-down philosophy that can be utilized to analyze FMs in any domain: model scope, model pretraining, and model design. Additionally, we systematically categorize PFM evaluation tasks into slide-level, patch-level, multimodal, and biological tasks, providing comprehensive benchmarking criteria. Our analysis identifies critical challenges in both PFM development (pathology-specific methodology, end-to-end pretraining, data-model scalability) and utilization (effective adaptation, model maintenance), paving the way for future directions in this promising field. Resources referenced in this survey are available at https://github.com/BearCleverProud/AwesomeWSI.
@@ 1 Introduction
Computational Pathology (CPath), the computational analysis of patient specimens (i.e., Whole Slide Images, WSIs), is increasingly important due to the critical role of histopathology. For gigapixel WSIs, Multiple Instance Learning (MIL) is the de facto framework, involving WSI patch partitioning, feature extraction via pretrained neural networks, and feature aggregation into WSI-level features Xiong et al. (2024b). Therefore, the MIL performance hinges on two components: the pretrained neural network (extractor) and the aggregator.
Pathology Foundation Models (PFMs), large-scale neural networks pretrained on extensive pathological data that learn generalizable features across diverse downstream tasks, such as HIPT Chen et al. (2022) and UNI Chen et al. (2024), mark a paradigm shift for MIL. Conventionally, due to lack of PFMs, ResNet-50 He et al. (2016) pretrained on ImageNet serves as the extractor Xiong et al. (2024a), but struggles with pathology-specific characteristics like minimal color variation, rotation-agnosticism, and hierarchical tissue organization. While limited labeled WSIs prevented supervised pretraining, Self-Supervised Learning (SSL) enables PFMs that exhibit superior generalizability in morphology recognition. This overcomes natural image pretraining limitations, where features mainly capture general visual attributes like edges and textures, enabling better performance on downstream pathology tasks even with limited data.
@@
Recent surveys for PFMs have made significant contributions to understanding PFMs Chanda et al. (2024), yet the field lacks a systematic taxonomy for technical analysis and a systematic organization of the evaluation tasks of PFMs. To address this critical gap, we present this survey with three primary contributions: 1) a hierarchical taxonomy organizing PFMs based on scope, training strategy, and design to enable holistic analysis, transferable to general vision FMs; 2) a comprehensive analysis of evaluation methodologies, examining their technical merits and limitations; and 3) a structured analysis of pathology-centric research challenges prioritizing underexplored directions. The manuscript is organized as follows: Section 2 formally formulates MIL and SSL; Section 3 introduces the proposed taxonomy; Section 4 examines evaluation tasks; Section 5 delineates future research directions; and Section 6 concludes the survey.
@@ 2.1 Multiple Instance Learning
A WSI is often represented as a bag of an unordered instances (patches). The goal of MIL is to predict the WSI label $\hat{Y}$ using only ground truth label $Y$ as supervision, without patch-level labels $\{y_{i}\}_{i=1}^{N}$, where $N$ is the number of patches in the bag. The relationship between the bag label $Y$ and instance labels $\{y_{i}\}_{i=1}^{N}$ is formally expressed as Xiong et al. (2023),
$Y=\begin{cases}1&\exists i,y_{i}=1,\\ 0&\forall i,y_{i}=0,\\ \end{cases}$ (1)
The implementation of MIL involves tessellating WSIs into non-overlapping patches $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times h\times w\times 3}$, with $h,w$ stand for height and width, respectively. These patches undergo feature extraction through an extractor $\mathcal{M}_{e}(\cdot)$, generating corresponding features $\boldsymbol{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d}$, where each feature is computed as $\boldsymbol{z}_{i}=\mathcal{M}_{e}(\boldsymbol{x}_{i})$, and $d$ is the hidden dimension of the extractor. Subsequently, an aggregator $\mathcal{M}_{g}(\cdot)$ agglomerates these features to form a bag-level feature $\boldsymbol{h}=\mathcal{M}_{g}({\boldsymbol{Z}})$ of the WSI, which serves as the input for the classification layer. Throughout aggregation, the extractor $\mathcal{M}_{e}(\cdot)$ usually remains frozen and non-trainable, while the aggregation network $\mathcal{M}_{g}(\cdot)$ is optimized during training.
@@ 2.2 Self-supervised Learning
SSL leverages unlabeled data by automatically generating supervisory signals through pretext tasks Ericsson et al. (2022). Given an input image $\boldsymbol{x}$, a transformation function $\mathcal{T}(\cdot)$ is applied to generate a modified version $\tilde{\boldsymbol{x}}=\mathcal{T}(\boldsymbol{x})$ and a corresponding pseudo-label $\tilde{y}$. An extractor $\mathcal{M}_{e}(\cdot)$ extracts features from $\tilde{\boldsymbol{x}}$ and generates a predicted label $\hat{y}=\mathcal{M}_{e}(\tilde{\boldsymbol{x}})$. The learning objective can be formalized as minimizing the difference between the predicted label $\hat{y}$ and the pseudo-label $\tilde{y}$. Common pretext tasks include contrastive learning, self-distillation, masked image modeling, etc., each designed to force the model to learn meaningful semantic features of the data. Through this process, the extractor can learn transferable features for downstream tasks on massive unlabeled data.
@@ 3 Hierarchical Taxonomy
Our taxonomy systematically organizes PFMs through three interdependent dimensions and reflect a top-down design philosophy: 1) Model Scope: a categorization of the scope of the PFMs, differentiating between PFMs focused on extractors, aggregators, and both components; 2) Model Pretraining: a detailed examination of the spectrum of image-centric pretraining methods, including slide-level, patch-level, and multimodal techniques; and 3) Model Design: a rigorous analysis of architecture, categorizing PFMs according to their number of parameters and scale. This top-down structure enables systematic comparisons of PFMs, as demonstrated in Table 1.
@@ 3.1 Model Scope
Extractor-centric models constitute the predominant approach in PFM development, driven by two factors: the importance of high-quality features and the necessity to address domain shift brought by ImageNet-pretrained CNNs. The role of the extractor aligns with established clinical practice, where pathologists emphasize cellular morphological analysis at the patch level. CTransPath Wang et al. (2022) pioneering the extractor training with a hybrid CNN-Transformer designs through Semantic-Relevant Contrastive Learning (SRCL) on 15 million patches. REMEDIS Azizi et al. (2023) demonstrates the feature extraction capability of ResNet-50 is constraint by domain shift across different medical imaging domains. Various advancements, including Virchow Vorontsov et al. (2024) and SINAI Campanella et al. (2024), further stress the significance of robust extractors.
@@
Aggregator-centric models plays a vital role in slide-level tasks as they are the only trainable models under direct supervision of ground truth labels, yet they are relatively underexplored compared to the extractor. CHIEF Wang et al. (2024), leveraging supervised pretraining with anatomical site to create an anatomy-aware aggregator, first demonstrates the efficacy of aggregator pretraining. More recent research like MADELEINE Jaume et al. (2025), TITAN Ding et al. (2024), and THREAD Vaidya et al. (2025), utilize multimodal data in aggregator pretraining with frozen patch features to enhance performance across downstream tasks. This paradigm shift reflects growing awareness that the aggregator critically impacts downstream task performance, particularly in low-resource clinical scenarios Xu et al. (2024a). This observation aligns with transfer learning principles, wherein pretraining on large-scale datasets effectively alleviates downstream data scarcity challenges. However, empirical evidence also reveals that the pretrained CHIEF aggregator occasionally performs worse than linear probing of the extractor Ding et al. (2024), which potentially attributable to the small model size when trained on pratraining-scale dataset, or the conflicts between domain bias and generic features. Consequently, further investigations are warranted to assess the advantages from pretrained larger aggregators.
@@
Hybrid-centric models are PFMs pretraining both the extractor and aggregator. Its advantage lies in full exploitation of the aggregator, as the aggregators can flexibly adapt to features from the extractor with pretraining-size data. HIPT pioneers this approach through hierarchical pretraining of the first two layers of extractor, excluding the last layer, which is substantiated through empirical performance. Similarly, Prov-GigaPath Xu et al. (2024a) pretrains a ViT extractor and a LongNet Ding et al. (2023) slide encoder; however, LongNet generates instance-level features rather than a single slide-level feature, necessitating integration of ABMIL Ilse et al. (2018) or non-parametric pooling strategies for slide-level tasks. mSTAR Xu et al. (2024b) distinguishes itself as the sole fully-pretrained hybrid-centric model by an inverted pretraining sequence: first optimizing the multimodal aggregator, followed by pretraining the extractor with the aggregator, which contrasts with the conventional paradigm.
@@
Analysis of recent developments reveals two observations. First, research emphasis has progressively shifted from feature extractor pretraining toward aggregator pretraining, a transition potentially attributable to both the robust performance of existing extractors and the increasing awareness of aggregator significance, especially in limited-data scenarios. Second, current aggregators demonstrate a hierarchical dependency pattern, wherein each successive model builds upon the capabilities of prior models. For instance, TITAN utilizes features from CONCHv1.5, which in turn leverages UNI as its encoder, thereby forming a cascading performance dependency chain where the efficacy of TITAN is inherently contingent upon CONCHv1.5 and, by extension, UNI, ultimately.
@@ 3.2 Model Pretraining
The pretraining methods can be categorized into supervised and SSL methods, with SSL prevailing due to their capabilities in capturing morphological patterns without labeled data, while only CHIEF opted for supervised pretraining for the aggregator. Based on our surveyed papers, SSL can be further divided into two main categories: vision-only and inter-modal methods. Vision-only methods employ three SSL techniques: contrastive learning (SimCLR, MoCov3), masked image modeling (MIM, MAE), and self-distillation (iBOT, DINO, DINOv2). In contrast, inter-modal methods often employ multi-stage pretraining, utilizing contrastive learning methods (CLIP, CoCa) for effective cross-modal alignment, before which, unimodal encoders are pretrained independently. We focus on methodology contributions in this section and present the details of each method, including input modalities and magnification and resolution of the patches, in the Model Pretraining column of Table 1.
@@
Contrastive Learning is an SSL branch that learns representations by maximizing similarity between positive pairs while minimizing that between negative pairs. Several seminal approaches have advanced this field: 1) SimCLR Chen et al. (2020) established foundational techniques such as aggressive data augmentation and large batch sizes; 2) MoCov3 Chen et al. (2021) advanced self-supervised learning for ViT through stabilized training techniques; 3) CLIP Radford et al. (2021) expanded the paradigm to multi-modal learning through large-scale image-caption pair training; and 4) CoCa Yu et al. (2022) proposed a unified method incorporating both contrastive and captioning objectives, enabling simultaneous visual-textual alignment and text generation capabilities. In medical domain, REMEDIS utilize SimCLR to enhance the robustness and data-efficiency in medical imaging. Pathoduet Hua et al. (2024) enhanced MoCov3 through the integration of cross-scale positioning and cross-stain transferring tasks, specifically addressing the challenges of stain transferability and tissue-level heterogeneity. CLIP is adapted for both extractors (PLIP Huang et al. (2023)) and aggregators (Prov-GiGapath, mSTAR, MADELEINE, and THREAD), due to its versatility in aligning two or more modalities. Notably, KEEP Zhou et al. (2024) has proposed a Knowledge-Enhanced Vision-Language (KEVL) pretraining, further adapting CLIP for extractor by incorporating domain expertise through knowledge-graph-cleaned image-text pairs. There are several applications of CoCa, both on the extractor and aggregator: CONCH Lu et al. (2024) adopts this framework to pretrain an extractor on 1.17 million image-caption pairs, enhancing both zero- and few-shot capabilities, while PRISM Shaikovski et al. (2024) and TITAN utilize CoCa to pretrain aggregators with multimodal capabilities.
@@
Masked Image Modeling is an SSL method that learns representations by predicting masked portions of images from their visible regions. SimMIM Xie et al. (2022) advanced the field by simplifying existing approaches through random masking and lightweight prediction head, and MAE He et al. (2022) introduced an asymmetric encoder-decoder design with high masking ratios. Recent investigations have demonstrated the efficacy of MIM in pretraining extractors; notably, SINAI Campanella et al. (2024) employs MAE to pretrain ViT models on an scale of 3.2 billion patches, establishing its scalability in pathological contexts. Similarly, MUSK Xiang et al. (2025) and BEPH Yang et al. (2024) further validate MIM by implementing BEiT-3 and BEiTv2 architectures, respectively. Additionally, Prov-GigaPath employs MAE to pretraing their slide encoder LongNet, demonstrating the efficacy of this method on aggregator pretraining.
@@
Self-distillation enables model learning through its own predictions across different views, simultaneously acting as teacher and student. DINO Caron et al. (2021) pioneered the use of self-distillation by employing a teacher-student architecture with momentum encoder and multi-crop training, while iBOT Zhou et al. (2022) performs MIM via self-distillation with an online tokenizer, and DINOv2 Oquab et al. (2023) refined the DINO framework through accelerating and stabilizing the training at scale. The efficacy of self-distillation for extractor has been demonstrated, as exemplified by several investigations: Phikon Filiot et al. (2023) implements iBOT on a corpus of 43 million patches spanning 16 distinct cancer sites; Phikon-v2 employs DINOv2 on 456 million patches derived from 30 cancer sites; RudolfV Dippel et al. (2024) incorporates DINOv2 with pathologist knowledge on 58 tissue types and 129 stains; and Hibou Nechaev et al. (2024) further extends DINOv2 on 1.2 billion patches. Additionally, the application of self-distillation extends beyond the extractor, as evidenced by TITAN Ding et al. (2024), which utilizes iBOT for general-purpose aggregator learning. Collectively, these investigations demonstrate the capacity and efficacy of self-distillation in PFM pretraining. In addition, there are methodological improvements customized for pathology in this category: 1) PLUTO Juyal et al. (2024) utilizes DINOv2 together with MAE objective and Fourier losses on 195 million patches; 2) GPFM Ma et al. (2024) proposes Unified Knowledge Distillation (UKD), incorporating MIM, self-distillation and expert knowledge distillation together as training objectives; 3) Virchow2 Zimmermann et al. (2024) enhances DINOv2 by applying pathology-specific augmentation and reducing tissue redundancy.
@@
Nature Wang et al. (2024) CHIEF P: Pretrained S: Sup.+CLIP P: CTransPath S: CHIEF Public + In-house 60,530 H&E WSIs $\sim$ 15 Million Patches [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]

### v2

@@ Abstract
Computational pathology, which involves analyzing whole slide images for automated cancer diagnosis, relies on multiple instance learning, where performance depends heavily on the feature extractor and aggregator. Recent Pathology Foundation Models (PFMs), pretrained on large-scale histopathology data, have significantly enhanced both the extractor and aggregator, but they lack a systematic analysis framework. In this survey, we present a hierarchical taxonomy organizing PFMs through a top-down philosophy applicable to foundation model analysis in any domain: model scope, model pretraining, and model design. Additionally, we systematically categorize PFM evaluation tasks into slide-level, patch-level, multimodal, and biological tasks, providing comprehensive benchmarking criteria. Our analysis identifies critical challenges in both PFM development (pathology-specific methodology, end-to-end pretraining, data-model scalability) and utilization (effective adaptation, model maintenance), paving the way for future directions in this promising field. Resources referenced in this survey are available at https://github.com/BearCleverProud/AwesomeWSI.
@@ 1 Introduction
Computational Pathology (CPath), the computational analysis of patient specimens (i.e., Whole Slide Images, WSIs), is increasingly important due to the critical role of histopathology. For gigapixel WSIs, Multiple Instance Learning (MIL) is the de facto framework, involving WSI patch partitioning, feature extraction via pretrained neural networks, and feature aggregation into WSI-level features Xiong et al. (2024b). Therefore, MIL performance hinges on two components: the pretrained neural network (extractor) and the aggregator.
Pathology Foundation Models (PFMs), neural networks pretrained on extensive pathological data that can be directly leveraged for diverse downstream tasks without retraining, such as HIPT Chen et al. (2022) and UNI Chen et al. (2024), mark a paradigm shift for MIL. Conventionally, due to the lack of PFMs, ResNet-50 He et al. (2016) pretrained on ImageNet Deng et al. (2009) serves as the extractor Xiong et al. (2024a), but struggles with pathology-specific characteristics like minimal color variation, rotation-agnosticism, and hierarchical tissue organization. While limited labeled WSIs prevented supervised pretraining, Self-Supervised Learning (SSL) enables PFMs that exhibit superior generalizability in morphology recognition. This overcomes natural image pretraining limitations, in which features mainly capture general visual attributes like edges and textures, enabling better performance on downstream tasks even with limited data.
@@
Recent surveys on PFMs have contributed significantly to the understanding of the field; however, these works either focus primarily on the impact of PFMs on the real world rather than technical investigations of them Ochi et al. (2025), or detail the previous efforts in this field without a systematic taxonomy for technical analysis and a systematic organization of the evaluation tasks of PFMs Chanda et al. (2024). To address these critical gaps, we introduce a comprehensive and timely survey of the current landscape. We collected papers from high-impact journals, including Nature, Nature Medicine, Nature Biomedical Engineering, Medical Image Analysis, as well as top-tier conferences such as CVPR, ICML, and AAAI. Given the rapid evolution of the field, we also incorporated preprints from repositories such as arXiv, bioRxiv, and medRxiv, acknowledging that many influential works are still under review. In total, our survey includes 27 PFM papers, 12 of which are preprints that have not yet been accepted by peer-reviewed conferences or journals.
We present this survey with three primary contributions: 1) a hierarchical taxonomy organizing PFMs based on scope, training strategy, and design to enable holistic analysis, transferable to general vision FMs; 2) a comprehensive analysis of evaluation methodologies, examining their technical merits and limitations; and 3) a structured analysis of pathology-centric research challenges prioritizing underexplored directions. The manuscript is organized as follows: Section 2 formally formulates MIL and SSL; Section 3 introduces the proposed hierarchical taxonomy; Section 4 examines evaluation tasks for PFMs; Section 5 delineates future research directions in this field; and Section 6 concludes our survey.
@@ 2.1 Multiple Instance Learning
In the MIL framework, a WSI is typically represented as a bag of $N$ unordered instances (or patches). The central objective of MIL is to predict the WSI-level label $\hat{Y}$ using only the ground truth bag label $Y$ as supervision, without access to the ground truth instance-level labels $\{y_{i}\}_{i=1}^{N}$. This setting reflects a common scenario in computational pathology, where obtaining slide-level annotations is feasible, but annotating individual patches is prohibitively expensive and impractical. The relationship between the bag label $Y$ and the instance labels $\{y_{i}\}_{i=1}^{N}$ is typically defined under standard MIL assumptions such as the presence-based assumption, and can be formally expressed as Xiong et al. (2023),
$Y=\begin{cases}1&\exists i,y_{i}=1\\ 0&\forall i,y_{i}=0\\ \end{cases}.$ (1)
The implementation of MIL involves tessellating WSIs into non-overlapping patches $\boldsymbol{X}=\{\boldsymbol{x}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times h\times w\times 3}$, with $h,w$ standing for height and width, respectively. These patches undergo feature extraction through an extractor $\mathcal{M}_{e}(\cdot)$, generating corresponding features $\boldsymbol{Z}=\{\boldsymbol{z}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d}$, where each feature is computed as $\boldsymbol{z}_{i}=\mathcal{M}_{e}(\boldsymbol{x}_{i})$, and $d$ is the hidden dimension of the extractor. Subsequently, an aggregator $\mathcal{M}_{g}(\cdot)$ agglomerates these features to form a bag-level feature $\boldsymbol{h}=\mathcal{M}_{g}({\boldsymbol{Z}})$ of the WSI, which finally serves as the input for the classification layer. Throughout aggregation, the extractor $\mathcal{M}_{e}(\cdot)$ usually remains frozen and non-trainable due to GPU memory constraints, while the aggregation network $\mathcal{M}_{g}(\cdot)$ is optimized during training. We refer readers unfamiliar with MIL to prior surveys for more details Carbonneau et al. (2018); Waqas et al. (2024).
@@ 2.2 Self-supervised Learning
SSL leverages unlabeled data by automatically generating supervisory signals through pretext tasks Ericsson et al. (2022). Given an input image $\boldsymbol{x}$, a transformation function $\mathcal{T}(\cdot)$ is applied to generate a modified version $\tilde{\boldsymbol{x}}=\mathcal{T}(\boldsymbol{x})$ and a corresponding pseudo-label $\tilde{y}$. An extractor $\mathcal{M}_{e}(\cdot)$ extracts features from $\tilde{\boldsymbol{x}}$ and generates a predicted label $\hat{y}=\mathcal{M}_{e}(\tilde{\boldsymbol{x}})$. The learning objective can be formalized as minimizing the difference between the predicted label $\hat{y}$ and the pseudo-label $\tilde{y}$. Common pretext tasks include contrastive learning, self-distillation, masked image modeling, etc., each designed to force the model to learn meaningful semantic features of the data. Through this process, the extractor can learn transferable features for downstream tasks on massive unlabeled data. We refer readers who are unfamiliar with SSL to prior surveys for more details Ericsson et al. (2022); Shurrab and Duwairi (2022); Gui et al. (2024).
@@ 3 Hierarchical Taxonomy
Our taxonomy systematically organizes PFMs through three interdependent dimensions and reflects a top-down design philosophy: 1) Model Scope: a categorization of the scope of the PFMs, differentiating between PFMs focused on extractors, aggregators, and both components; 2) Model Pretraining: a detailed examination of the spectrum of image-centric pretraining methods, including slide-level, patch-level, and multimodal techniques; and 3) Model Design: a rigorous analysis of architecture, categorizing PFMs according to their number of parameters and scale. This top-down structure enables systematic comparisons of PFMs, as shown in Table 1.
@@ 3.1 Model Scope
Extractor-centric models constitute the predominant approach in PFM development, driven by two factors: the importance of high-quality features and the necessity to address domain shift brought by ImageNet-pretrained CNNs. The role of the extractor aligns with established clinical practice, where pathologists emphasize cellular morphological analysis at the patch level. CTransPath Wang et al. (2022) pioneers the extractor training with a hybrid CNN-Transformer design through Semantic-Relevant Contrastive Learning (SRCL) on 15 million patches. REMEDIS Azizi et al. (2023) demonstrates that the feature extraction capability of ResNet-50 is constrained by domain shift across different medical imaging domains. Various advancements, including Virchow Vorontsov et al. (2024) and SINAI Campanella et al. (2024), further stress the significance of robust extractors.
@@
Aggregator-centric models play a vital role in slide-level tasks as they are the only trainable models under direct supervision of ground truth labels, yet they are relatively underexplored compared to the extractor. CHIEF Wang et al. (2024b), leveraging supervised pretraining with the anatomical site to create an anatomy-aware aggregator, first demonstrates the efficacy of aggregator pretraining. More recent research like MADELEINE Jaume et al. (2025), TITAN Ding et al. (2024), and THREAD Vaidya et al. (2025) utilizes multimodal data in aggregator pretraining with frozen patch features to enhance performance across downstream tasks. This paradigm shift reflects growing awareness that the aggregator critically impacts downstream task performance, particularly in low-resource clinical scenarios Xu et al. (2024a). This observation aligns with transfer learning principles, wherein pretraining on large-scale datasets effectively alleviates downstream data scarcity challenges. However, empirical evidence also reveals that the pretrained CHIEF aggregator occasionally performs worse than linear probing of the extractor Ding et al. (2024), which is potentially attributable to the small model size when trained on a pretraining-scale dataset, or to the conflicts between domain bias and generic features. Consequently, further investigations are warranted to assess the advantages of pretrained larger aggregators.
@@
Hybrid-centric models are PFMs that pretrain both the extractor and aggregator. Their advantage lies in full exploitation of the aggregator, as the aggregators can flexibly adapt to the extractor with pretraining-scale data. HIPT pioneers this approach through hierarchical pretraining of the first two layers of the extractor, excluding the last layer, which is substantiated through empirical performance. Similarly, Prov-GigaPath Xu et al. (2024a) pretrains a ViT extractor and a LongNet Ding et al. (2023) slide encoder; however, LongNet generates instance-level features rather than a single slide-level feature, necessitating integration of ABMIL Ilse et al. (2018) or non-parametric pooling strategies for slide-level tasks. TANGLE Jaume et al. (2024) pretrains both a ViT feature extractor and a transcriptomics-guided ABMIL aggregator. Finally, mSTAR Xu et al. (2024b) distinguishes itself as a fully-pretrained hybrid-centric model by an inverted pretraining sequence, contrasting with the conventional paradigm: first optimizing the multimodal aggregator, followed by pretraining the extractor with the aggregator.
@@
Analysis of recent developments reveals two observations. First, research emphasis has progressively shifted from feature extractor pretraining toward aggregator pretraining, a transition potentially attributable to both the robust performance of existing extractors and the increasing awareness of aggregator significance, especially in limited-data scenarios. Second, current aggregators demonstrate a hierarchical dependency pattern, wherein each successive model builds upon the capabilities of prior models. For instance, TITAN utilizes features from CONCHv1.5, which in turn leverages UNI as its encoder, thereby forming a cascading performance dependency chain where the efficacy of TITAN is inherently contingent upon CONCHv1.5 and, by extension, UNI.
@@ 3.2 Model Pretraining
The pretraining methods can be categorized into supervised and SSL methods, with SSL prevailing due to their capabilities in capturing morphological patterns without labeled data, while only CHIEF opted for supervised pretraining for the aggregator. Based on our surveyed papers, SSL can be further divided into two main categories: vision-only and inter-modal methods. Vision-only methods employ three SSL techniques: contrastive learning (SimCLR, MoCov3), masked image modeling (MIM, MAE), and self-distillation (iBOT, DINO, DINOv2). In contrast, inter-modal methods often employ multi-stage pretraining, utilizing contrastive learning methods (CLIP, CoCa) for effective cross-modal alignment before which unimodal encoders are pretrained independently. We focus on methodology contributions in this section and present the details of each method, including input modalities, magnification, and resolution of the patches, in the Model Pretraining column of Table 1.
@@
Contrastive Learning is an SSL branch that learns representations by maximizing similarity between positive pairs while minimizing that between negative pairs. Several seminal approaches have advanced this field: 1) SimCLR Chen et al. (2020) established foundational techniques such as aggressive data augmentation and large batch sizes; 2) MoCov3 Chen et al. (2021) advanced self-supervised learning for ViT through stabilized training techniques; 3) CLIP Radford et al. (2021) expanded the paradigm to multi-modal learning through large-scale image-caption pair training; and 4) CoCa Yu et al. (2022) proposed a unified method incorporating both contrastive and captioning objectives, enabling simultaneous visual-textual alignment and text generation capabilities. In the medical domain, REMEDIS utilizes SimCLR to enhance the robustness and data efficiency in medical imaging. TANGLE adopts a revised SimCLR method with gene expression reconstruction and slide subset alignment. Pathoduet Hua et al. (2024) enhanced MoCov3 through the integration of cross-scale positioning and cross-stain transferring tasks, specifically addressing the challenges of stain transferability and tissue-level heterogeneity. CLIP is adapted for both extractors (PLIP Huang et al. (2023)) and aggregators (Prov-GiGapath, mSTAR, MADELEINE, and THREAD), due to its versatility in aligning two or more modalities. Notably, KEEP Zhou et al. (2024) has proposed a Knowledge-Enhanced Vision-Language (KEVL) pretraining, further adapting CLIP for the extractor by incorporating domain expertise through knowledge-graph-cleaned image-text pairs. There are several applications of CoCa, both on the extractor and aggregator: CONCH Lu et al. (2024) adopts this framework to pretrain an extractor on 1.17 million image-caption pairs, enhancing both zero- and few-shot capabilities, while PRISM Shaikovski et al. (2024) and TITAN utilize CoCa to pretrain aggregators with multimodal capabilities.
@@
Masked Image Modeling is an SSL method that learns representations by predicting masked portions of images from their visible regions. SimMIM Xie et al. (2022) advanced the field by simplifying existing approaches through random masking and a lightweight prediction head, and MAE He et al. (2022) introduced an asymmetric encoder-decoder design with high masking ratios. Recent investigations have demonstrated the efficacy of MIM in pretraining extractors; notably, SINAI Campanella et al. (2024) employs MAE to pretrain ViT models on a scale of 3.2 billion patches, establishing its scalability in pathological contexts. Similarly, MUSK Xiang et al. (2025) and BEPH Yang et al. (2024) further validate MIM by implementing BEiT-3 and BEiTv2 architectures, respectively. Additionally, Prov-GigaPath employs MAE to pretrain its slide encoder LongNet, demonstrating the efficacy of this method on aggregator pretraining.
@@
Self-distillation enables model learning through its own predictions across different views, simultaneously acting as teacher and student. DINO Caron et al. (2021) pioneered the use of self-distillation by employing a teacher-student architecture with momentum encoder and multi-crop training, while iBOT Zhou et al. (2022) performs MIM via self-distillation with an online tokenizer, and DINOv2 Oquab et al. (2023) refined the DINO framework by accelerating and stabilizing the training at scale. The efficacy of self-distillation for the extractor has been demonstrated by several investigations: Phikon Filiot et al. (2023) implements iBOT on a corpus of 43 million patches spanning 16 distinct cancer sites; Phikon-v2 Filiot et al. (2024) employs DINOv2 on 456 million patches derived from 30 cancer sites; RudolfV Dippel et al. (2024) incorporates DINOv2 with pathologist knowledge on 58 tissue types and 129 stains; and Hibou Nechaev et al. (2024) further extends DINOv2 on 1.2 billion patches. Additionally, the application of self-distillation extends beyond the extractor, as evidenced by TITAN Ding et al. (2024), which utilizes iBOT for general-purpose aggregator learning. These investigations demonstrate the capacity of self-distillation in PFM pretraining. In addition, there are methodological improvements customized for pathology in this category: 1) PLUTO Juyal et al. (2024) utilizes DINOv2 together with MAE objective and Fourier losses on 195 million patches; 2) GPFM Ma et al. (2024) proposes Unified Knowledge Distillation (UKD), incorporating MIM, self-distillation and expert knowledge distillation together as training objectives; 3) Virchow2 Zimmermann et al. (2024) enhances DINOv2 by applying pathology-specific augmentation and reducing tissue redundancy.
@@
Nature Wang et al. (2024b) CHIEF P: Pretrained S: Sup.+CLIP P: CTransPath S: CHIEF Public + In-house 60,530 H&E WSIs $\sim$ 15 Million Patches [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
CVPR Jaume et al. (2024) TANGLE P: iBOT S: Alignment P: ViT-B/16 S: ABMIL TG-GATEs 47,227 WSIs 6,597 Image-Gene Pair [Uncaptioned image] [Uncaptioned image]